# Étale motives of geometric origin

### v1

@@ Abstract
The theorem holds when $X$ is the spectrum of a field $k$ and $\Lambda$ be a ring which is flat over the integers and regular.
@@
Hence, the dual of $C$ is also a torsion motive.
Writing $C=\rho_{!}D$, then $D$ is a dualisable object of $\mathrm{D}(k_{\mathrm{\acute{e}t}},\Lambda[1/p])$. In particular, it belongs to the thick stable subcategory generated by the $f_{\sharp}\underline{\Lambda[1/p]}$ for $f$ finite étale according to [Rui22, Proposition 2.1.2.15]. Thus, its image $C$ through $\rho_{!}$ is geometric as $\rho_{!}$ is compatible with the functors $f_{\sharp}$ and $f^{*}$ for $f$ finite étale using [CD16, §4.4.2]. Finally, $M\oplus M[1]$ is geometric and therefore $M$ is geometric. ∎
Proposition 4.4.
4.1 holds if the ring $\Lambda$ is flat over the integers and regular.
@@
Using 3.2 and 3.8, we can assume $X$ to be noetherian. This is now an adaptation of the arguments given in [CD16, Lemma 6.3.25]. Let $M$ be a constructible motive on $X$. By noetherian induction, using the localisation property and the stability of geometric motives under pushforwards by closed immersions it suffices to show that there is an open immersion $j\colon U\to X$ such that $j^{*}(M)$ is geometric, i.e. we can replace $X$ with any non-empty open subscheme. Therefore, we can assume $X$ to be irreducible. Let $\eta$ be the generic point of $X$. Consider the commutative diagram:
@@
Then, 4.3 above implies that $\beta$ is an equivalence and 3.2 implies that $\alpha$ is an equivalence. On the other hand $u$ is fully faithful and $v$ is an equivalence by 3.8. Hence, they are both equivalences. The fact that $u$ is an equivalence yields the result. ∎
The following lemma was considered and proven by Cisinski and Déglise with a particular class of rings.
Lemma 4.5.
Let $\Lambda$ be a commutative ring and let $X$ be a noetherian scheme and let $M\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$, then there exist an étale covering $U$ of $X$ such that the restriction $M_{\mid U}$ is geometric.
Proof.
The proof is merely an improvement of [CD16, Theorem 6.3.26]. By dévissage and localisation, it is enough to prove the claim for $M$ dualisable. Let $x$ be a point of $X$, and let $S=\mathrm{Spec}\ \mathscr{O}_{X,x}^{\mathrm{sh}}$ be the strict henselisation of $X$ at $x$. The map $u\colon S\to X$ is the limit of the maps $v\colon V\to X$ for $V$ ranging the étale neighborhoods of $x$ in $X$. For $\mathcal{C}$ a symmetric monoidal $\infty$-category, denote by $\mathcal{C}^{\mathrm{db}}$ the full subcategory of dualisable objects. The functor
$\mathrm{colim}_{V}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\Lambda)^{\mathrm{db}}\to\mathrm{DM}^{\mathrm{\acute{e}t}}(S,\Lambda)$
is fully faithful. Indeed, given $M,N\in\mathrm{DM}^{\mathrm{\acute{e}t}}(V_{0},\Lambda)$, dualisable, we can compute the following mapping spectra:
$\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\Lambda)}(v^{*}M,v^{*}N)=\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\mathbb{Z})}(\mathbb{Z}_{V},v^{*}(M^{\vee}\otimes_{\Lambda}N)),$
and the same over $S$ if one replaces $v$ by $u$. The result then follows from [CD16, Remark 6.3.8], but we give the arguments for the reader’s convenience. Being an equivalence can be checked after tensoring by $\mathbb{Q}$ and $\mathbb{Z}/p\mathbb{Z}$. The rationalised part is an equivalence because by 3.1 it is identified with
$\mathrm{colim}_{v}\hskip 2.84526pt\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\mathbb{Q})}(\mathbb{Q}_{V},v^{*}((M^{\vee}\otimes_{\Lambda}N)\otimes_{\mathbb{Z}}\mathbb{Q}))\to\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(S,\mathbb{Q})}(\mathbb{Q}_{S},u^{*}((M^{\vee}\otimes_{\Lambda}N)\otimes_{\mathbb{Z}}\mathbb{Q})),$
which is an equivalence because continuity always holds for the presentable $\infty$-categories and $\mathbb{Q}$ is a compact object of $\mathrm{DM}^{\mathrm{\acute{e}t}}(V_{0},\mathbb{Q})$. The part tensored by $\mathbb{Z}/p\mathbb{Z}$ is an equivalence by rigidity: we are computing étale cohomology in a colimit of topoï, and this holds thanks to [CD16, Lemma 1.1.12 and Theorem 1.1.5].
In particular when applied to the generic points $\eta$ of $X$, as $S=\overline{\eta}$ in this case by 3.7 the image of a dualisable object in $\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\Lambda)$ in $\mathrm{DM}^{\mathrm{\acute{e}t}}(\overline{\eta},\Lambda)$ is constructible, hence geometric, thus the functor
$\mathrm{colim}_{V}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}(V,\Lambda)^{\mathrm{db}}\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(\overline{\eta},\Lambda)$
is an equivalence, and there by the usual properties of spreading out there exists an étale neighborhood $V_{\eta}$ of $\eta$ such that our dualisable object $M$ of $X$ is geometric when restricted to $V_{\eta}$. As $X$ is noetherian, there is an étale covering $U$ of $X$ such that $M_{\mid U}$ is geometric as an object of $\mathrm{DM}^{\mathrm{\acute{e}t}}(U,\Lambda)$. ∎
Proposition 4.6.
Let $\Lambda$ be a ring of finite type over the integers and let $X$ be a finite-dimensional noetherian scheme. Then 4.1 holds in $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$.
Proof.
By assumption we can write $\Lambda=R/I$ with $R$ a finite type polynomial ring over the integer. By 4.1 above the result holds over $R$. Let $M\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$. By 4.5 there exist an étale covering $j:U\to X$ such that $j^{*}M$ is geometric. In particular, $j^{*}M$ is a direct factor of a finite colimit of objects of the form $N\otimes_{R}\Lambda$ with $N\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,R)=\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,R)$. As $R/I$ is perfect over $R$, each $N\otimes_{R}\Lambda$ is in fact constructible as objects of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,R)$, so that by étale descent $M\otimes_{R}\Lambda$ is a constructible $R$-motive, hence geometric, and $M$ is a direct factor of $M\otimes_{R}\Lambda$. ∎
From this we deduce the theorem for slightly more general coefficients.
Corollary 4.7.
Let $\Lambda$ be a ring of finite type over the integers. Then 4.1 holds for $\Lambda$.
Proof.
Assume $X=\lim X_{i}$ is the limit of $X_{i}$ finite-dimensional noetherian schemes. By 4.6 for each $X_{i}$ we have an equivalence
@@
$\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X_{i},\Lambda)\simeq\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X_{i},\Lambda).$
In particular, the canonical functor
is fully faithful. The result is therefore equivalent to the essential surjectivity of this functor. For this we may use 3.9: there exists a geometric object $P\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda)$ and a map $P\to M\oplus M[1]$ with a cofiber $C$ such that there is a $n\in\mathbb{N}^{*}$ such that $C/n\simeq C\oplus C[1]$. Thus it suffices to know that the constructible object $C/n$ is in the image. As $C/n$ is in the image of the forgetful functor
$\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda),$
by 4.8 it suffices to prove that $C/n$ is in the image of the functor
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{D}_{\mathrm{cons}}(X_{i},\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})\to\mathrm{D}_{\mathrm{cons}}(X,\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}).$
Because over noetherian schemes we have proven that constructible and geometric motives coincide with coefficient $\Lambda$, modding out by $(n)$ the mapping spectra gives the full faithfulness of the above functor. The claim is that it is essentially surjective. For this, an easy dévissage reduces to the case of a dualisable object $K$. Let $j:U\to X$ an finite étale covering such that $j^{*}K$ is constant perfect, this exists thanks to [MW22, Corollary 7.4.12]. Let $P$ be the value of this constant sheaf on $U$. Then $P$, seen as a $\Lambda$-module, is perfect: it can be written as a finite colimit of free $\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}$ modules, and $\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}$ has a finite resolution by finite free $\Lambda$-modules. As $P\otimes_{\Lambda}(\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})$ has $P$ as a direct factor, we can assume that our $P$ is of the for $P^{\prime}/n$ for some perfect complex $P^{\prime}\in\mathrm{Perf}_{\Lambda}$. In particular, the cosimplicial object $(j^{m}_{\sharp}\underline{P^{\prime}}_{U_{m}})$ in $\mathrm{D}_{\mathrm{cons}}(X,\Lambda)$ associated to the Čech nerve of $j$ is uniformly bounded, so that its totalisation is finite. This gives that $K=\lim_{\Delta}j^{m}_{\sharp}P^{\prime}/n$ as a finite limit. Now $j$, hence all the $j^{m}$, are defined over some $X_{i}$, so that it is clear that $K$ is pulled back from some $X_{i}$. ∎
The following lemma was used in the proof of the previous corollary.
Lemma 4.8.
Let $\Lambda$ be a derived commutative $\mathbb{Z}/n\mathbb{Z}$-algebra. Then the natural functor
$\mathrm{D}(X_{\mathrm{\acute{e}t}},\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$
is an equivalence, which restricts to $\mathrm{D}_{\mathrm{cons}}(X_{\mathrm{\acute{e}t}},\Lambda)\simeq\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$. Here the definitions one has to take for the objects used in this lemma are exactly the same as the one used everywhere else in this paper.
Proof.
The equivalence of presentable $\infty$-categories is formal: this is true for $\Lambda=\mathbb{Z}/n\mathbb{Z}$, and then we obtain the one we considered by tensoring by $\mathrm{Mod}_{\Lambda}$ over $\mathrm{Mod}_{\mathbb{Z}/n\mathbb{Z}}$. The fact that it restrict to categories of constructible objects follows from the fact that both definition are strictly the same. ∎
Proof.
(of 4.1) We write $\Lambda=\mathrm{colim}_{i}\hskip 2.84526pt\Lambda_{i}$ with $\Lambda_{i}$ a ring of finite type over the integers. For each ring $\Lambda_{i}$ we know that $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda_{i})=\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda_{i})$. Thus the functor
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda_{i})\to\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$
is fully faithful an the theorem would follow from the fact that it is essentially surjective. By 3.9 as before we reduce to proving that an object of the form $C/n$, for $C\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$, is in the image, but then this reduces to proving that $C/n$ is in the image of the fully faithful functor
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{D}_{\mathrm{cons}}(X,\Lambda_{i}\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})\to\mathrm{D}_{\mathrm{cons}}(X,\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}).$
The above functor is indeed fully faithful because for each $i$ we know that
$\mathrm{D}_{\mathrm{cons}}(X,\Lambda_{i}\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})=\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda_{i}\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})$
and $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda_{i})=\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda_{i})$, so the full faithfulness follows from continuity of geometric motives, mod $n$. Now to prove that the functor is essentially surjective, we may assume that the object $K$ we are trying to reach is dualisable, and even of the form $j_{\sharp}P$ with $P$ perfect constant and $j\colon U\to X$tale, as in the proof of 4.7. By [HRS23, Lemma 2.2] there is a $P_{i}$ perfect over $\Lambda_{i}$ such that $P=P_{i}\otimes_{\Lambda_{i}}\Lambda$, and the proof is finished. ∎
Remark 4.9.
Let $X$ be a qcqs finite-dimensional scheme and let $\Lambda$ be a ring of $n$ torsion, with $n$ an integer which is invertible on $X$. Then, by rigidity, the content of 4.1 is that the category $\mathrm{D}_{\mathrm{cons}}(X,\Lambda)$ is generated by the $f_{\sharp}(\Lambda_{Y})(n)$ for $f\colon Y\to X$ smooth and $n$ an integer. This result fails in the $\ell$-adic setting over a finite field because not all $\ell$-adic complexes are mixed in the sense of [BBD82, Section 5].
@@
Since the presentable $\infty$-category $\mathrm{DM}^{\mathrm{\acute{e}t}}(-,\Lambda)$ satisfies Nisnevich descent, the localisation property and proper base change, it satisfies cdh-descent as in the proof of [CD19, Theorem 3.3.10] applies in this level of generality. Thus $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies non effective cdh descent. The effectivity of descent follows from cdh-excision: a descent data of geometric motives gives us a motive as a finite limit of proper pushforwards, which is itself geometric by 3.5 (1). For hyperdescent, this follows from the fact that the cdh topos is hypercomplete over qcqs schemes of finite valuative dimension by [EHIK21a, Corollary 2.4.16]. ∎
@@
Let $\Lambda$ be a commutative ring. The functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies h-descent in the sense of [Stacks, Tag 0ETQ] over qcqs finite-dimensional schemes.
@@
By writing $S=\lim_{i}S_{i}$ where the $S_{i}$ are affine $\mathbb{Z}$-schemes of finite type, we can find an index $i_{0}$ such that $f$ is the pullback of a finite map $f_{i_{0}}\colon T_{i_{0}}\to S_{i_{0}}$ using [Gro66, Théorème 8.10.5]. By noetherian induction, we can find a finite stratification $\{S_{i_{0}}^{1},\ldots,S_{i_{0}}^{p}\}$ of $S_{i_{0}}$ such that $S_{i_{0}}^{k}$ is open in $\bigcup_{k^{\prime}\geqslant k}S_{i_{0}}^{k^{\prime}}$ for each $k$ and $f_{i_{0}}$ is the composition of a finite étale map and a purely inseparable map over each $S_{i_{0}}^{k}$. Hence, pulling back this stratification to $S$, we get a stratification $\{S^{1},\ldots,S^{p}\}$ of $S$ with the same properties.
@@
thus the left and right vertical maps of the above diagram are equivalences by assumption, so that the middle one also. Therefore, by induction on the number $p$ of strata of $S$, this proves the full faithfulness by induction, given that we have descent along finite étale maps by 2.4 and that localisation ensures that $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ is an equivalence along purely inseparable maps.
@@
Assume first that $\Lambda$ is a polynomial ring over the integers. Then we proceed exactly as in the proof of [EHIK21b, Theorem 5.6]: The functor
@@
has a right adjoint given by the formula
$\mathrm{fib}(f_{*}f^{*}M\oplus i_{*}i^{*}M\to f_{*}k_{*}k^{*}f^{*}M),$
because all pushforwards preserve geometric motives by [CD16, Theorem 6.2.13]. By localization, the family of functors $(f^{*},i^{*})$ is conservative, thus the functor 5.3.1 is also conservative. Hence given geometric motives $M_{Y}$, $M_{Z}$ and $M_{W}$ over $Y$, $Z$ and $W$ such that we have equivalences $k^{*}(M_{Y})\simeq M_{W}\simeq g^{*}(M_{Z})$, it suffices to show that the maps $f^{*}(M)\to M_{Y}$ and $i^{*}(M)\to M_{Z}$ are equivalences, with $M$ the cofiber of the morphism $f_{*}(M_{Y})\oplus i_{*}(M_{Z})\to f_{*}k_{*}(E_{W})$, that is, the right adjoint of 5.3.1 is fully faithful. Using 3.1, we can therefore assume $\Lambda$ to be a $\mathbb{Q}$-algebra or to be a $\mathbb{Z}/p\mathbb{Z}$-algebra for some prime number $p$. The first case follows from [EHIK21b, Theorem 5.6, Lemma 2.1] while the second case follows from [BM21, Theorem 5.14] and [EHIK21a, Corollary 3.2.11].
Now for a ring $\Lambda$ of finite type over the integers that we write as $R/I$ with $R$ a polynomial ring, we proceed as follows: first, the functor 5.3.1 is fully faithful. Indeed, it suffices to check that the unit of the adjunction is an equivalence when evaluated on the generators of geometric motives which are of the form $M\otimes_{R}(R/I)$ with $M\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,R)$. As $R/I$ is perfect over $R$ (because $R$ is noetherian) the operation $-\otimes_{R}(R/I)$ is only a finite colimit, whence the fully faithfulness. Now given descent data $(M_{Y},M_{Z},M_{W},isos)$ in the limit, the three motives $M_{Y}$, $M_{Z}$ and $M_{W}$ map to geometric objects of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,R)$, together with descent data because $\Lambda$ is perfect over $R$. Thus there exists a $N\in\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,R)$ whose restrictions to $Y$, $Z$ and $W$ are $M_{Y}$, $M_{Z}$ and $M_{W}$. Thus $N\otimes_{R}\Lambda$ is an object of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda)$ with an image in the fibre product of $\infty$-categories containing the data $(M_{Y},M_{Z},M_{W},isos)$ as a direct factor. The full faithfulness implies that $(M_{Y},M_{Z},M_{W},isos)$ is in the image. Now given an arbitrary commutative ring $\Lambda$, we may write it as a filtered colimit of finite type rings over the integers, and using 3.4 the result is therefore a consequence of the commutation of finite limits with filtered colimits. ∎

### v2

@@ Abstract
Let $\Lambda=\mathrm{colim}\Lambda_{i}$ be a filtered colimit of rings and let $X$ be a Noetherian scheme. Then the natural functor
$\mathrm{colim}_{i}\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda_{i})\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$
is an equivalence.
Proof.
This proof was communicated to the authors by Denis-Charles Cisinski. Fully faithfulness can be checked after tensorisation by $\mathbb{Q}$ and $\mathbb{Z}/p\mathbb{Z}$. The rationalised statement follows from 3.8 and the fact that $\Lambda\mapsto\mathrm{Mod}_{\Lambda}(\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\mathbb{Z}))$ commutes with filtered colimits ([HA, Theorem 4.8.5.11]). The mod $p$ statement can be proven as follows: a dévissage reduces to showing the statement for dualisable objects, and then for maps out of the constant sheaf, and even from maps out of the constant sheaf $\mathbb{Z}$, and it is now a consequence of commutativity of étale cohomology with uniformly bounded below complexes as in [HRS23, Lemma 5.3].
Now for the essential surjectivity first using 3.9, as it is clear that geometric objects are in the image, we reduce to reaching an object of the form $C/n$ which is of the form $\rho_{!}K$ with $K$ in $\mathrm{D}_{\mathrm{cons}}(X,\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})$. Then, a dévissage reduces to a dualisable object for which by 3.3 there exists an étale covering $f\colon U\to X$ such that $f^{*}(K)$ is constant with value a complex $P$ in $\mathrm{Perf}_{\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}}$. As in the proof of 3.4 we see that $K=\mathrm{colim}_{n\in\Delta}(f_{n})_{\sharp}P$ with $(f_{n})$ the Čech nerve of $f$. As $P$ is of the form $Q\otimes_{\Lambda_{i}\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}}(\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z})$ (this follows from the fact that the objects of $\mathrm{Perf}_{(-)}$ are the compact objects of $\mathrm{Mod}_{(-)}$ which commutes with colimts), we see that $K=(\mathrm{colim}_{n\in\Delta}(f_{n})_{\sharp}Q)\otimes_{\Lambda_{i}\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}}\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}$ is the image of a dualisable object. ∎
Lemma 4.4.
Let $\Lambda$ be a commutative ind-regular ring. The theorem holds when $X$ is the spectrum of a field $k$.
@@
Hence, the dual of $C$ is also a torsion motive. It suffices to show that any such $C$ is geometric.
Assume first that $\Lambda$ is a regular ring. Writing $C=\rho_{!}D$, then $D$ is a dualisable object of $\mathrm{D}(k_{\mathrm{\acute{e}t}},\Lambda[1/p])$. In particular, it belongs to the thick stable subcategory generated by the $f_{\sharp}\underline{\Lambda[1/p]}$ for $f$ finite étale according to [Rui22, Proposition 2.1.2.15]. Thus, its image $C$ through $\rho_{!}$ is geometric as $\rho_{!}$ is compatible with the functors $f_{\sharp}$ and $f^{*}$ for $f$ finite étale using [CD16, §4.4.2].
Now we let $\Lambda$ be an ind-regular commutative ring, we write it as $\Lambda=\mathrm{colim}_{i}\Lambda_{i}$ with each $\Lambda_{i}$ regular. By 4.3 any object of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(k,\Lambda)$ comes from a object in $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(k,\Lambda_{i})$ for some $i$, hence is geometric by the above paragraph. ∎
@@
(of 4.1) Using 3.5 and 3.10, we can assume $X$ to be Noetherian. This is now an adaptation of the arguments given in [CD16, Lemma 6.3.25]. Let $M$ be a constructible motive on $X$. By Noetherian induction, using the localisation property and the stability of geometric motives under pushforwards by closed immersions it suffices to show that there is an open immersion $j\colon U\to X$ such that $j^{*}(M)$ is geometric, i.e. we can replace $X$ with any non-empty open subscheme. Therefore, we can assume $X$ to be irreducible. Let $\eta$ be the generic point of $X$. Consider the commutative diagram:
@@
Then, 4.4 above implies that $\beta$ is an equivalence and 3.5 implies that $\alpha$ is an equivalence. On the other hand $u$ is fully faithful and $v$ is an equivalence by 3.10. Hence, they are both equivalences. The fact that $u$ is an equivalence yields the result. ∎
Example 4.5.
This result fails without the assumption that $\Lambda$ is regular: take $\Lambda=\mathbb{F}_{2}[x]/(x^{2}-1)$. Then, by rigidity, we have
@@
$\mathrm{DM}^{\mathrm{\acute{e}t}}(\mathrm{Spec}(\mathbb{R}),\Lambda)\underset{\rho_{!}}{\xleftarrow{\sim}}\mathrm{Sh}(\mathbb{R}_{\mathrm{\acute{e}t}},\Lambda)\simeq\mathrm{D}(\Lambda[C_{2}])$
where $C_{2}$ is the cyclic group with $2$ elements. Taking dualisable objects, we get an equivalence between $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(\mathrm{Spec}(\mathbb{R}),\Lambda)$ and the subcategory $\mathrm{D}_{\Lambda\text{-perf}}(\Lambda[C_{2}])$ of $\mathrm{D}(\Lambda[C_{2}])$ made of those complexes with a perfect underlying complex of $\Lambda$-modules.
If $f\colon X\to\mathbb{R}$ is smooth, this equivalence sends $M_{\mathbb{R}}(X)(n)$ to the complex $f_{\sharp}(\Lambda)(n)$. The latter is isomorphic to $(f_{\sharp}(\mathbb{F}_{2})(n))\otimes_{\mathbb{F}_{2}}\Lambda$. As by [BG23, Corollary 4.5], $f_{\sharp}(\mathbb{F}_{2})(n)$ belongs to the thick subcategory generated by $\mathbb{F}_{2}$ and $\mathbb{F}_{2}[C_{2}]$, the complex $f_{\sharp}(\Lambda)(n)$ therefore belongs to the thick subcategory $\mathrm{D}_{\mathrm{perm}}(\Lambda[C_{2}])$ generated by $\Lambda$ and $\Lambda[C_{2}]$. Hence, through this equivalence $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(\mathrm{Spec}(\mathbb{R}),\Lambda)$ corresponds to $\mathrm{D}_{\mathrm{perm}}(\Lambda[C_{2}])$.
Consider the complex $\Lambda_{x}$ of $\Lambda[C_{2}]$-modules whose underlying complex is $\Lambda$ placed in degree 00 and with the action of the generator of $C_{2}$ given by multiplication by $x$. The complex $\Lambda_{x}$ belongs to $\mathrm{D}_{\Lambda\text{-perf}}(\Lambda[C_{2}])$ but by [BG22, Example 3.15] it does not belong to $\mathrm{D}_{\mathrm{perm}}(\Lambda[C_{2}])$. Hence, the inclusion $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(\mathrm{Spec}(\mathbb{R}),\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(\mathrm{Spec}(\mathbb{R}),\Lambda)$ is strict.
In fact, the result fails even for a characteristic 00 non-regular ring: take $\Lambda^{\prime}=\mathbb{Z}[x]/(x^{2}-1)$, and consider the $\Lambda^{\prime}[C_{2}]$-module $\Lambda^{\prime}_{x}$, defined in the same way. It can be seen as an étale motive through the Artin motive functor $\rho_{!}\colon\mathrm{Sh}(\mathbb{R}_{\mathrm{\acute{e}t}},\Lambda^{\prime})\to\mathrm{DM}^{\mathrm{\acute{e}t}}(\mathrm{Spec}(\mathbb{R}),\Lambda^{\prime})$. This object is dualisable, and if it were geometric then its image under the functor $-\otimes_{\Lambda^{\prime}}\Lambda$, which is $\Lambda_{x}$ would also be geometric, which is not.
@@
Since the presentable $\infty$-category $\mathrm{DM}^{\mathrm{\acute{e}t}}(-,\Lambda)$ satisfies Nisnevich descent, the localisation property and proper base change, it satisfies cdh-descent as the proof of [CD19, Theorem 3.3.10] applies in this level of generality. Thus $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies non effective cdh descent. The effectivity of descent follows from cdh-excision: a descent data of geometric motives gives us a motive as a finite limit of proper pushforwards, which is itself geometric by 3.6 (1). For hyperdescent, this follows from the fact that the cdh topos is hypercomplete over qcqs schemes of finite valuative dimension by [EHIK21a, Corollary 2.4.16]. ∎
@@
Let $\Lambda$ be a commutative ind-regular ring. The functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies h-descent in the sense of [Stacks, Tag 0ETQ] over qcqs finite-dimensional schemes.
@@
By writing $S=\lim_{i}S_{i}$ where the $S_{i}$ are affine $\mathbb{Z}$-schemes of finite type, we can find an index $i_{0}$ such that $f$ is the pullback of a finite map $f_{i_{0}}\colon T_{i_{0}}\to S_{i_{0}}$ using [EGA IV, Théorème 8.10.5]. By Noetherian induction, we can find a finite stratification $\{S_{i_{0}}^{1},\ldots,S_{i_{0}}^{p}\}$ of $S_{i_{0}}$ such that $S_{i_{0}}^{k}$ is open in $\bigcup_{k^{\prime}\geqslant k}S_{i_{0}}^{k^{\prime}}$ for each $k$ and $f_{i_{0}}$ is the composition of a finite étale map and a purely inseparable map over each $S_{i_{0}}^{k}$. Hence, pulling back this stratification to $S$, we get a stratification $\{S^{1},\ldots,S^{p}\}$ of $S$ with the same properties.
@@
thus the left and right vertical maps of the above diagram are equivalences by assumption, so that the middle one also. Therefore, by induction on the number $p$ of strata of $S$, this proves the full faithfulness by induction, given that we have descent along finite étale maps by 2.5 and that localisation ensures that $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ is an equivalence along purely inseparable maps.
@@
Because pushforwards preserve geometric motives over quasi-excellent schemes by [CD16, Theorem 6.2.13], the functor
@@
admits a right adjoint. By localization, the family of functors $(f^{*},i^{*})$ is conservative thus the above functor is conservative. To show that it is an equivalence, it suffices to show that its right adjoint is fully faithful. Thus take geometric motives $M_{Y}$, $M_{Z}$ and $M_{W}$ over $Y$, $Z$ and $W$ such that we have equivalences $k^{*}(M_{Y})\simeq M_{W}\simeq g^{*}(M_{Z})$. Let $M$ be the fiber of the map $f_{*}(M_{Y})\oplus i_{*}(M_{Z})\to f_{*}k_{*}(E_{W})$. It is the image under the right adjoint of 5.3.1 of the object $(M_{Y},M_{Z},W)$. To finish the proof, it suffices to show that the natural maps $f^{*}(M)\to M_{Y}$ and $i^{*}(M)\to M_{Z}$ are equivalences. Assume first that the ring $\Lambda$ is flat over the integers. Using 3.1, we can therefore assume $\Lambda$ to be a $\mathbb{Q}$-algebra or to be a $\mathbb{Z}/p\mathbb{Z}$-algebra for some prime number $p$. The first case follows from [EHIK21b, Theorem 5.6, Lemma 2.1] while the second case follows from [BM21, Theorem 5.14] and [EHIK21a, Corollary 3.2.11]. Now for a ring $\Lambda$ of finite type over the integers that we write as $R/I$ with $R$ a polynomial ring, we proceed as follows: as $\Lambda$ is perfect over $R$, the three motives $M_{Y}$, $M_{Z}$ and $M_{W}$ map to geometric objects of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,R)$, together with descent data. In particular, the maps $f^{*}(M)\to M_{Y}$ and $i^{*}(M)\to M_{Z}$ are equivalences in $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,R)$ hence in $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$. Thus the functor 5.3.1 is an equivalence whenever $\Lambda$ is of finite type over the integers. Finally for a general ring, we can write it as a colimit of ring of finite type over the integers and with 4.3 this is just a consequence of the commutation of filtered colimits with finite limits. ∎